OpenStudy (anonymous):

if the measure of each interior angle of a regular polygon is 176 find the number of sides on the polygon. Show all work.

OpenStudy (anonymous):

my bad, it's the formula for vertex angle. .

OpenStudy (anonymous):

what is the formula of a vertex angle?

OpenStudy (anonymous):

 where n is the number of sides. .

OpenStudy (experimentx):

(n-2)x(180)/n = 176 --- solve this equation for 'n', 'n' is the number of sides

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (anonymous):

how about using this. .

OpenStudy (anonymous):

given that the sum of the interior angles of a triangle s a180, and to find the measure of the vertex angle it use the formula 360/n where n is pertains to the number of sides, in a regular polygon, from the center, if we create a triangle, it is always isosceles, sometimes equilateral,. use this.. measure of the 1 interior angle of on the base =(sum of the interior angle - vertex angle)/2

OpenStudy (anonymous):

176 = 180 - 360/n n =90 i got 90 also

OpenStudy (anonymous):

ooh, i got it even without the aid of the interior angle formula. . Yepey! lol

OpenStudy (anonymous):

haha. But I know this is bad but I kind of cheated and the answer book says it's 40

OpenStudy (anonymous):

huh? lol

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

yes hey tell me this would it be 40 if the number is 171?

OpenStudy (anonymous):

if it's 171, but the given is 176.. :-)

OpenStudy (anonymous):

no the given being 171